DiVA and KoF11

Kurt Otto 2012-02-07

\underline{Why DiVA?}

That the IT department should be well represented in DiVA with good order among the items serves several purposes:

1. To provide a reliable and reasonably complete database of our publications for research colleagues

2. Convenient generation of publication lists on our web pages

3. To ascertain appropriate faculty funding, which will partly be determined by the number of refereed publications in DiVA

4. Damage control in bibliometric studies

[figure]

My work in DiVA

involves (but is not limited to)

• collection of new publication items either through instructions to the authors or by entering them myself;

• correction of new items from the authors or those imported by the University Library (UL), especially regarding research program, subject category, and CAS-id;

• maintenance of the guidelines (internt/web/diva) especially at “årsredovisning”;

• cleaning backwards in time (when there is time left).

Hunting ISI-id in February 2011

As a precaution in the data collection for the bibliometrics in KoF11, I examined 193 DiVA items (entered or “cleaned up” by me) from IT (2007–2010). The result by March 8th, 2011, was that

• 87 were not in ISI Web of Science

• 34 had been provided with ISI-id by UL (their job as promised)

• 72 I had to provide with ISI-id

My opinions on bibliometrics

Generally,

• it is a dubious attempt to measure the impact of research

• it is potentially harmful when used for distributing research funds (especially in the hands of lazy administrators)

Especially for IT (in KoF11),

• it is inaccurate, since the (estimated) CI coverage is only 58% and the data collection is (rather) incomplete, v.s.

\underline{What to do?}

Be prepared and be aware!